# ChroniclingAmericaQA: A Large-scale Question Answering Dataset based on Historical American Newspaper Pages

### v1

@@ 4.4. Human Evaluation
To assess the quality of the ChroniclingAmericaQA dataset, we conducted a manual evaluation study. We randomly selected 360 question-answer pairs for evaluation, sampling 30 pairs from each decade represented in the dataset. Six graduate students were asked to rate these pairs on a scale of 1 to 5, with 1 being ”very bad” and 5 being ”very good.” The evaluators were instructed to assess the pairs based on four criteria: the readability of the question, ensuring it is grammatically correct and flows well; the readability of the answer, verifying its grammatical correctness and fluency; relevance, to ascertain if the generated question pertains to the provided passage; and non-ambiguity, to ensure the question is straightforward and clear.
Each criterion’s human evaluation results were averaged and presented in Table 7. The evaluations yielded high scores across all metrics, indicating that our dataset is of high quality, particularly regarding the relevance and readability of both questions and answers. The score for non-ambiguity was also notably high, demonstrating that most questions are direct and unambiguous.
@@ 5. Use cases
Our dataset can be used in several ways. First, it provides a new benchmark for training and evaluating QA and MRC models on historical texts, enabling models to handle the complexities of historical text, including OCR inaccuracies and language evaluation. The evaluation can be done for different spans of time within the time frame of the data (1800-1920). It can be also conditioned on different regions of USA or particular states.
Second, since our dataset contains scanned images of news papers as well as the corresponding raw and cleaned OCR texts, it is possible to benchmark models on each of these. Some memory institutions may still hold only scans of historical materials, while others, such as in the case of Chronicling America, may contain quite noisy OCRed text. Our dataset allows then benchmarking models on these different types of inputs, offering realistic scenarios in which the models may need to be required to operate.
Third, the dataset can be used to increase the public’s engagement with historical documents. For example, it can help people build skills in critical reading and language arts.
Finally, ChroniclingAmericaQA could potentially serve as an educational resource; for example, it could be used by teachers and educators to assess students’ reading comprehension on rare, historical materials. Note, that our dataset generation framework can be applied on other historical news collections.
@@ 6. Conclusion
In this work, we introduce ChroniclingAmetricaQA, a large-scale question-answering dataset comprising 485k question-answer pairs over a collection of historical American newspapers with the objective of facilitating the development of QA and MRC systems over historical texts. Our dataset is unique as it is specifically based on historical newspapers covering the longest time period of 120 years from 1800 to 1920 years, among other available datasets on news documents. We highlight the difficulties of utilizing historical newspapers for QA tasks, such as the poor quality of OCR text and the need for post-processing to enhance text quality. We employed language models such as GPT 3.5 Turbo for OCR text correction and the T5-base model for question generation to mitigate these challenges. We also carried out a comprehensive evaluation of our dataset on various transformer models and LLMs, demonstrating the effectiveness of our dataset. Models trained on ChroniclingAmericaQA consistently surpass the performance of those fine-tuned on other datasets.
Ethical considerations Lastly, we would like to mention ethical considerations of our work. Since the dataset has been created from temporally distant data, there is possibility that some questions may be offensive or biased in relation to some ethnic groups or other segments of society. We made sure that the possibility of such questions in our dataset is low by performing post-processing search using a set of prepared keywords coupled with careful manual analysis. Nevertheless, this kind of issue is not specific to our dataset but is a problem typical in many heritage materials (Borenstein et al., 2023), including also ones in the Chronicling America repository, and should be continuously studied and mitigated in the future.

### v2

@@ 4.4. Human Evaluation
To assess the quality of the ChroniclingAmericaQA dataset, we also conducted a manual evaluation study. We randomly selected 360 question-answer pairs for evaluation, sampling 30 pairs from each decade represented in the dataset. Six graduate students were asked to rate these pairs on a scale of 1 to 5, with 1 being ”very bad” and 5 being ”very good.” The evaluators were instructed to assess the pairs based on four criteria: the readability of the question, ensuring it is grammatically correct and flows well; the readability of the answer, verifying its grammatical correctness and fluency; relevance, to ascertain if the generated question pertains to the provided passage; and non-ambiguity, to ensure the question is straightforward and clear. Each question was annotated by one of the annotators.
Each criterion’s results were averaged and are presented in Table 7. The evaluations gave high scores across all metrics, indicating that our dataset is of good quality, particularly regarding the relevance and readability of both questions and answers. The score for non-ambiguity was also notably high, demonstrating that most questions are direct and unambiguous.
@@ 5. Use cases
Our dataset can be used in several ways. First, it provides a new benchmark for training and evaluating QA and MRC models on historical texts, enabling models to handle the complexities of historical documents, including OCR inaccuracies and language evolution. The evaluation can be done for arbitrary, different spans of time within the time frame of the dataset (1800-1920). It can be also conditioned on different regions of USA or particular states.
Second, since our dataset contains scanned images of news papers as well as the corresponding raw and cleaned OCR texts, it is possible to benchmark models on each of these. Some memory institutions still hold only scans of historical materials, while others, such as in the case of the Chronicling America, may contain quite noisy OCRed text. Our dataset allows then benchmarking models on these different types of inputs, offering realistic scenarios in which the models may need to be required to operate.
Third, our dataset can be utilized in a similar manner to datasets such as SQuAD (Rajpurkar et al., 2016), ArchivalQA (Wang et al., 2022), and StreamingQA (Liska et al., 2022) for contributing to the ongoing efforts in enhancing the adaptability and knowledge retention of large language models (Hu et al., 2023; Tack et al., 2024).
Finally, ChroniclingAmericaQA could potentially serve as an educational resource or as a tool to increase the public’s engagement with historical documents; for example, it could be used by teachers and educators to assess students’ reading comprehension on rare, historical materials. Note, that our dataset generation framework can be applied on other historical news collections.
@@ 6. Conclusion
In this work, we introduce ChroniclingAmetricaQA, a large-scale question-answering dataset comprising 487k question-answer pairs over a collection of historical American newspapers with the objective of facilitating the development of QA and MRC systems over historical texts. Our dataset is unique as it is specifically based on historical newspapers covering the longest time period of 120 years (from 1800 to 1920) among available datasets based on news documents. We highlight the difficulties of utilizing historical newspapers for QA tasks, such as the poor quality of OCR text and the need for post-processing to enhance text quality. We employed language models such as GPT 3.5 Turbo for OCR text correction and the T5-base model for question generation to mitigate these challenges. We also carried out a comprehensive evaluation of our dataset on various transformer models and LLMs, demonstrating the effectiveness of our dataset and providing reference baselines. Models trained on ChroniclingAmericaQA were found to consistently surpass the performance of those fine-tuned on other datasets.
Ethical considerations Lastly, we would like to mention ethical considerations of our work. Since the dataset has been created from temporally distant data, there is possibility that some questions may be offensive or biased in relation to some ethnic groups or other segments of society. We made sure that the possibility of such questions occurring in our dataset is low by performing post-processing search using a set of prepared keywords coupled with careful manual analysis. This kind of issue is not specific to our dataset but is a problem typical in many heritage materials (Borenstein et al., 2023), including also ones in the Chronicling America repository, and should be continuously studied and mitigated in the future.